a pickup truck accelerates 3.25m/s2 from rest, for 5.80 s. find it's final speed.


1
Expert's answer
2022-06-28T08:19:06-0400

The final velocity

"v_f=v_i+at\\
\rm =0+3.25*5.80=18.9\:\rm m/s"